OpenStudy (anonymous):

Debate:Mechanical Advantage Give real life examples of effort/load where Mechanical Advantage is present.

OpenStudy (anonymous):

MA=Load/Effort

OpenStudy (anonymous):

examples; used in simple machines like wheel and axle, levers,pulleys, and inclined planes

OpenStudy (anonymous):

@ganeshie8 @mathway Join the debate/discussion

OpenStudy (anonymous):

@Astrophysics

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (anonymous):

@Kaelyn78 @keshaun455

OpenStudy (anonymous):

@Robert136

OpenStudy (anonymous):

@danielasumo @Australopithecus @AmyRoseRules @Donblue @water_boy @whiterose109

OpenStudy (anonymous):

Solid start to the discussionXD

OpenStudy (anonymous):

Now waiting for invitees to join.

OpenStudy (anonymous):

@ashking1

OpenStudy (anonymous):

@Robert136

OpenStudy (anonymous):

let's start the debate

OpenStudy (anonymous):

in 5

OpenStudy (anonymous):

4

OpenStudy (anonymous):

3

OpenStudy (anonymous):

2

OpenStudy (anonymous):

1

OpenStudy (anonymous):

debate start

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

okay, let's start the debate off

OpenStudy (nincompoop):

there is nothing to debate about

OpenStudy (anonymous):

okay, how about discuss?

OpenStudy (nincompoop):

if you want to set up a debate, you need a proposition/assertion and one will support it and another will provide the negative

OpenStudy (anonymous):

okay, I'll start one tomorrow..thank you for the tips.. it's getting kind of late

OpenStudy (nincompoop):

cool! :)

OpenStudy (anonymous):

I'll tag you when I start it.@nincompoop

OpenStudy (astrophysics):

`Give me a lever long enough and a fulcrum on which to place it, and I shall move the world.` - Archimedes

OpenStudy (nincompoop):

I bet Archimedes was not joking, but I also bet that he wouldn't be able to

OpenStudy (astrophysics):

Lets calculate it

OpenStudy (astrophysics):

All we use is \[L_1 = \frac{ m_2 \times L_2 }{ m_1 }\] where L2 is the length of the lever, where we note archimedes himself is doing it, so he's 68 kg approximately, and the fulcrum being placed 16000 km from the earth. m2 is the mass of earth which is 5.972 x 10^(24) kg \[L_1 = \frac{ (5.972 \times 10^{24}kg) (16000 km) }{ (68kg) }\]

OpenStudy (astrophysics):

So the lever has to be 1.405x10^(27) km long xD

OpenStudy (anonymous):

you really weren't joking :) that's so cool

OpenStudy (astrophysics):

Haha nope, but good luck finding a lever that long

OpenStudy (jack1):

anything with gears lever or pulley system